OpenStudy (anonymous):

\[\log_{5} \log_{3} \log_{2^x} =0 then x is?

OpenStudy (anonymous):

\[\log_{5} \log_{3} \log_{2^x} =0 then x is?

OpenStudy (anonymous):

it is log x to the base 2.......

OpenStudy (anonymous):

log5 log3 log x to the base 2 =0 then x is ?

OpenStudy (anonymous):

plzz answer this with steps......

OpenStudy (mertsj):

Is it really the product of those logs?

OpenStudy (mertsj):

\[\log_{2}5(\log_{2}3)(\log_{2}x)=0 \]

OpenStudy (mertsj):

Is that the problem?

OpenStudy (anonymous):

log5 log3 log x to the base 2 =0 then x is ?

OpenStudy (anonymous):

log5 log3 log2^x

OpenStudy (anonymous):

this is the problem

OpenStudy (anonymous):

log2^x ( base 2)

OpenStudy (mertsj):

I am here. And I am not sure what the problem is. I still don't know if it is the product of the three logs or what. What class are you taking? Are you sure there are not plus or minus signs between those logs? Is there more to the problem such as the logs of 5 and 3 are given?

OpenStudy (anonymous):

it is the product i will take a photo

OpenStudy (mertsj):

ok

OpenStudy (mertsj):

Please include the directions with the picture.

OpenStudy (anonymous):

log5 log 3 log2^x =0 find x?

OpenStudy (anonymous):

\[\log_{5} \log_{3} \log_{2^{x}} =0 find \] x

OpenStudy (anonymous):

is it clear

OpenStudy (mertsj):

So it's not a product.

OpenStudy (mertsj):

It's the log of the log of the log. And remember that logs are exponents. And since you have a list of answers there, I would plug them in and see which one works.

OpenStudy (kropot72):

Is 1 included in the answers?

OpenStudy (anonymous):

log5 log 3 log2^x =0 find x? the question is to find x how it came log of log of log.....................the ans=8 plzz solve for it

OpenStudy (mertsj):

For example if x is 8 then you have \[\log_{2}8=3 \]

OpenStudy (anonymous):

yes u r right

OpenStudy (mertsj):

And the problem becomes: \[\log_{5}\log_{3}3 \]

OpenStudy (anonymous):